I have two cell phone plans. first one is $209.00 plus a monthly fee of $34.00. The second is $89.00 plus a monthly fee of $42.00. At what month is the cost for each plan the same?
M=number of months using the cell plan, and C= total cost for the cell plan.

Answer by mananth(16949)

I have two cell phone plans. first one is $209.00 plus a monthly fee of $34.00. The second is $89.00 plus a monthly fee of $42.00. At what month is the cost for each plan the same?
M=number of months using the cell plan, and C= total cost for the cell plan.
..
frst plan
C= 34M+209
Second Plan
C= 42M+89
..
Both have to be same.
34M+209=42M+89
42M-34M=209-89
8M=120
M=15